The work done to push a steel box to a distance of 485 cm is 118.825 joules. a) How much force is applied? b) What is the work done if the force act at θ = 40 ᵒ? C) if steel box is pushed in 35 sec, what is the power in kilowatts and horsepower ?


1
Expert's answer
2021-05-09T15:21:32-0400

a)


"F=\frac{W}{d}=\frac{118.825}{4.85}=24.5\ N"

b)


"W'=118.825\cos{40}=91.025\ J"

c)


"P=\frac{91.025}{35}=2.6\ W"